# Deconstructing the Goldilocks Zone of Neural Network Initialization

### v1

@@ Abstract
The second-order properties of the training loss have a massive impact on the optimization dynamics of deep learning models. Fort & Scherlis (2019) discovered that a high positive curvature and local convexity of the loss Hessian are associated with highly trainable initial points located in a region coined the “Goldilocks zone”. Only a handful of subsequent studies touched upon this relationship, so it remains largely unexplained. In this paper, we present a rigorous and comprehensive analysis of the Goldilocks zone for homogeneous neural networks. In particular, we derive the fundamental condition resulting in non-zero positive curvature of the loss Hessian and argue that it is only incidentally related to the initialization norm, contrary to prior beliefs. Further, we relate high positive curvature to model confidence, low initial loss, and a previously unknown type of vanishing cross-entropy loss gradient. To understand the importance of positive curvature for trainability of deep networks, we optimize both fully-connected and convolutional architectures outside the Goldilocks zone and analyze the emergent behaviors. We find that strong model performance is not necessarily aligned with the Goldilocks zone, which questions the practical significance of this concept.
@@ 1 Introduction
Figure 1: The Goldilocks zone is an area of high positive curvature of the Hessian. Left: Originally, the Goldilocks zone is observed for a narrow range of initialization scales. Middle: Setting the appropriate softmax temperature $T$ allows for high positive curvature at initialization of any norm. Right: Recreating the Goldilocks zone at an unscaled initialization just by varying $T$.
@@
While the Hessian is extensively studied throughout training and at convergence, fewer works focus on the initialization stage. Recently, Fort & Scherlis (2019) discovered the Goldilocks zone—a region of the optimization space marked by abnormally high values of positive curvature $\text{Tr}(H)/\lVert H\rVert_{F}$ and local convexity of the loss Hessian $H$ (see Figure 1). The anomalous readings of these metrics are recorded at a certain distance from the origin of the configuration space where some widely used initialization schemes such as Xavier (Glorot & Bengio, 2010) and Kaiming (He et al., 2015) are found. Thus, the Goldilocks zone is believed to be a hollow centered spherical shell that contains a high density of suitable initial points. Indeed, an appropriate parameter norm is crucial to avoid exploding and vanishing signals, and it seems reasonable that initializations of extreme norm (outside the Goldilocks zone) might suffer from this notorious issue. However, internal covariate shift is largely solved in practice by BatchNorm (Ioffe & Szegedy, 2015), and it can be directly accounted for in some special cases. For example, given an $L$-layer homogeneous network $f_{\alpha\theta}$ with an $\alpha$-scaled initialization (Dinh et al., 2017), the appropriate logit variance and gradient norm can be restored by applying a carefully selected softmax temperature $T=\alpha^{L}$ and learning rate $\eta=\mathcal{O}(\alpha^{2})$. These adjustments ensure that $f_{\alpha\theta}$ follows exactly the same training trajectory and has the same initial positive curvature as $f_{\theta}$ (Figure 1). Thus, contrary to the original claims by Fort & Scherlis (2019), the Goldilocks zone cannot be characterized by the initialization norm alone and is not even a subset of the configuration space. Instead, we find a more fundamental condition governing the positive curvature of the Hessian in Section 3.
@@
Previous works also attempted to describe the Goldilocks zone more formally. Fort & Ganguli (2019) develop an abstraction, the random logit model, in which they show that the positive curvature of cross-entropy loss vanishes with increased logit variance. Cohen et al. (2021) attribute the decrease in cross-entropy sharpness during optimization to the collapse of the softmax distribution, which also has implications for the Goldilocks zone. However, the phenomenon of positive curvature—its relation to the initialization norm and trainability—remains unexplained. In this work, focusing on homogeneous networks, we derive a comprehensive description of the Goldilocks zone from the Gauss-Newton decomposition (Section 3) and formally associate high positive curvature with certain properties of the network (Section 4). In particular, we show that near-zero curvature results from a relatively low spectral norm of the G-term in the Hessian decomposition due to either saturated softmax or vanishing logit gradients, which naturally arise for high- and low-norm initializations, respectively. Inside the Goldilocks zone, we prove that the highest positive curvature is observed for networks with low confidence, which in turn is associated with low initial loss and, for balanced datasets, with vanishing expected loss gradient.
@@
When the conditions governing the positive curvature of the loss are established, we inquire about their relation to model trainability and optimization dynamics. Fort & Scherlis (2019) find that optimization on random subspaces of the parameter space converges only if they intersect the Goldilocks zone. For unconstrained optimization, Gur-Ari et al. (2018) discovered that gradients are mainly confined to a low-rank subspace spanned by the Hessian top eigenvectors and vanish along flatter directions. Since the excess of positive curvature manifests a larger separation between bulk and outlier eigenvalues, it should then be associated with a more robust top-eigenspace and, intuitively, a more informative training signal. Motivated to unveil this relationship, we use gradient descent to optimize homogeneous networks using a wide spectrum of initialization norms and learning rates and taxonomize the emergent behaviors. Interestingly, we find that successful training is not necessarily correlated with the Goldilocks zone. We demonstrate setups where the slightest increase in the initialization norm of LeNet-5 leads to degenerate learning despite happening well within boundaries of the Goldilocks zone. These dynamics are marked by an increasing amount of zero logits and, to the best of our knowledge, we are the first to report this behavior.
@@
This paper conducts an extensive study of the Goldilocks zone of homogeneous neural networks, both analytically and empirically.
  • In Section 3, we demonstrate that the Goldilocks zone cannot be characterized by the initialization norm alone, as was previously claimed by Fort & Scherlis (2019). Instead, we derive a more fundamental condition resulting in non-zero positive curvature and find that it disappears due to saturated softmax on one end and vanishing logit gradients on the other.
  • In Section 4, we closely study the interior of the Goldilocks zone and analytically associate high positive curvature with low model confidence, low initial loss, and low cross-entropy gradient norm.
  • In Section 5, we report the evolution and performance of scaled homogeneous networks when optimized by gradient descent both inside and outside the Goldilocks zone. Our investigation shows that positive curvature is a poor estimator of the initialization trainability and exhibits a range of interesting effects for initializations on the edge.
All of our experiments are implemented in PyTorch (Paszke et al., 2017) and run on an internal cluster with NVIDIA RTX8000 and A100 GPUs. We use an open-source implementation of stochastic power iteration with deflation to compute the top eigenvalues and eigenvectors of the Hessian matrix (Golmant et al., 2018). Our code is available here: github.com/avysogorets/goldilocks-zone.
@@ 2 Preliminaries & Notation
By default, $T=1$, in which case we refer to softmax simply as $\sigma$. The corresponding cross-entropy loss is $\ell(p,y)=-\log p_{y}$. We restrict our analysis to homogeneous models satisfying $f_{\alpha\theta}=\alpha^{L}f_{\theta}$ for any scalar $\alpha>0$ where $L$ is the number of layers (weight matrices) in $f$. This is a rather technical assumption: homogeneous models are widely used in practice and include ReLU networks without biases. The Hessian matrix $H\in\mathbb{R}^{P\times P}$ holds the second derivatives of the loss at $\theta$: $H_{ij}=\partial^{2}\ell/\partial\theta_{i}\partial\theta_{j}$. In principle, all of the above quantities depend on one particular or a batch of inputs $(X^{\mu},y^{\mu})$; we use the superscript $\mu$ to make this dependence explicit where needed.
Positive curvature.
@@
In their seminal work, Fort & Scherlis (2019) define the Goldilocks zone rather informally as a region of high (non-zero) positive curvature of the loss Hessian, which can be unmistakably identified in Figure 1. Positive curvature of the loss Hessian is defined by
$\frac{\text{Tr}(H)}{\lVert H\rVert_{F}}=\frac{\sum_{i=1}^{P}\lambda_{i}}{\sqrt{\sum_{i=1}^{P}\lambda_{i}^{2}}}\,,$ (2)
@@
where $\{\lambda_{i}\}_{i=1}^{P}$ are the eigenvalues of $H$. Another metric used by Fort & Scherlis (2019) is local convexity, which refers to the fraction of positive eigenvalues. These two metrics are intimately related and can detect the Goldilocks zone independently of each other, so we focus only on positive curvature throughout this paper. Since computing the full Hessian is intractable for most modern architectures, Fort & Scherlis (2019) substitute it with a projection $H_{d}=R^{T}HR$ onto a low-rank random subspace with basis defined by a sparse matrix $R\in\mathbb{R}^{P\times d}$ with orthogonal columns. Given an initialization $\theta$, this amounts to computing the Hessian of the training loss of the model parameterized by $R\hat{\theta}+\theta$ with respect to latent $d$-dimensional parameters $\hat{\theta}$ at the origin of the chosen low-rank space, which is much more accessible. We adopt the same strategy and assume that the first- and second-order derivatives with respect to $\theta$ more generally represent derivatives with respect to trainable $d$-dimensional parameters, which can be latent parameters $\hat{\theta}$ or the original model weights if we let $d=P$ and $R=I_{P}$. This technical nuisance affects none of our analyses, but we add further comments as it becomes necessary.
@@ 3 Revisiting the Goldilocks Zone
Fort & Scherlis (2019) introduced the Goldilocks zone as a region of the parameter space with an excess of positive curvature and local convexity of the loss function, as defined in Section 2. Starting from a gold standard Kaiming initialization $\theta_{0}$, Fort & Scherlis (2019) record these statistics over a ray $\{\alpha\theta_{0}\colon\alpha>0\}$ and read unusually high values when $\alpha$ falls within a relatively narrow range centered around $\alpha=1$ (Figure 1 left). Moreover, they find that SGD constrained to a random subspace is successful only if it intersects the Goldilocks zone. On the basis of these observations, the authors suggest that the Goldilocks zone is a thick, hollow spherical shell about the origin in the configuration space, which is densely populated with initial points amenable for training.
A simple example shows that, strictly speaking, this visually appealing representation cannot be true. For homogeneous models, the $\alpha$-scale transformation $\theta\rightarrow\alpha\theta$ with $\alpha>0$ effectively scales the underlying configuration space by $\alpha$ (Dinh et al., 2017). We shall derive next that the cross-entropy loss landscape scales together with the configuration space when the softmax temperature satisfies $T=\alpha^{L}$, restoring the positive curvature of the scaled model $f_{\alpha\theta}$ to that of the original model $f_{\theta}$ (see the middle plot in Figure 1). Thus, we argue that initialization norm has a coincidental relationship to the Goldilocks zone, calling for an alternative, analytically driven characterization.
@@
$H_{ij}=\underbrace{\sum_{k=1}^{K}\sum_{c=1}^{K}\frac{\partial z_{k}}{\partial\theta_{i}}\left[\nabla^{2}_{z}\ell\right]_{kc}\frac{\partial z_{c}}{\partial\theta_{j}}}_{\text{G-term }(G_{*})}+\underbrace{\sum_{k=1}^{K}\left[\nabla_{z}\ell\right]_{k}\frac{\partial^{2}z_{k}}{\partial\theta_{j}\partial\theta_{i}}}_{\text{H-term }(H_{*})}\,,$ (4)
where $\nabla^{2}_{z}\ell=\text{diag}(p)-pp^{\top}$ is the Hessian of the loss with respect to the logits and $\nabla_{z}\ell=p-y$ where $p$ is the softmax output and $y$ is a one-hot encoded tagret (Cohen et al., 2021). To emphasize the dependence of the G-term and H-term on $p$ we will sometimes refer to them as $G_{*}(p)$ and $H_{*}(p)$, respectively. Note that $G_{*}(p)=J^{\top}[\text{diag}(p)-pp^{\top}]J$ where $J$ is the Jacobian matrix. For the $\alpha$-scaled model $f^{\prime}$ used together with softmax $\sigma_{T}$, we similarly get $\nabla^{2}_{z^{\prime}}\ell^{\prime}=T^{-2}\left[\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}\right]$ and $\nabla_{z^{\prime}}\ell^{\prime}=T^{-1}(p^{\prime}-y)$. Combining this with homogeneity of gradients, we derive the Gauss-Newton decomposition of the Hessian of $f^{\prime}$:
$H^{\prime}$ $=G^{\prime}_{*}+H^{\prime}_{*}=\frac{\alpha^{2L-2}}{T^{2}}G_{*}(p^{\prime})+\frac{\alpha^{L-2}}{T}H_{*}(p^{\prime}).$ (5)
Linearity of differentiation ensures that this equation is valid for the full-batch Hessian, too. Indeed, the full G-term and the full H-term are just the averages of the per-sample quantities $G^{\mu}_{*}$ and $H^{\mu}_{*}$ defined in Equation 4, respectively. Thus, we will abuse this notation and refer to their full counterparts in the same way where appropriate. In this case, the probability vectors $p$ and $p^{\prime}$ can be viewed as matrices.
Returning to our discussion on the shape of the Goldilocks zone, we remark that $T=\alpha^{L}$ yields $H^{\prime}=\alpha^{-2}H$ (as this temperature ensures $p=p^{\prime}$). Since the measures of positive curvature and local convexity are robust to scaling of the Hessian matrix, we conclude that initialization of any norm can be in the Goldilocks zone provided an appropriate softmax temperature (see Figure 1). Now that the connection between the initialization norm and the Goldilocks zone is compromised, we leverage Equation 5 to establish the more fundamental principles governing positive curvature in neural networks.
@@
The Gauss-Newton decomposition in Equation 4 is a common entry point for many studies on the Hessian of large neural networks. The Hessian exhibits a “bulk-outlier” eigenspectrum with the majority of eigenvalues small and clustered around zero and only a handful of large positive outliers (Sagun et al., 2016; Gur-Ari et al., 2018; Ghorbani et al., 2019). This decomposition is inherited from the individual spectra of $G_{*}$ and $H_{*}$ with the top and the bulk eigenvalues attributed to these two terms, respectively. We present a survey of works concerned with this phenomenon in Appendix A.
@@
We are now in position to describe the exact conditions that result in an excess of positive curvature—the hallmark of the Goldilocks zone. The eigenstructures of the individual terms in the Gauss-Newton decomposition of the Hessian suggest that the G-term is the one and only source of positive curvature of the Hessian, and so
$\lVert G^{\prime}_{*}\rVert_{2}\gtrsim\lVert H^{\prime}_{*}\rVert_{2}$ (6)
is a necessary and sufficient condition for the Goldilocks zone. We observe this exact correspondence in Figure 2. Otherwise, when $\lVert G^{\prime}_{*}\rVert_{2}$ is sufficiently small relative to $\lVert H^{\prime}_{*}\rVert_{2}$, the bulk-like eigenspectrum of the H-term swallows the outliers of the G-term and is inherited by the Hessian, resulting in a near-zero positive curvature. This sudden change in the eigenstructure is known as the BBP phase transition (Baik et al., 2005; Fort & Ganguli, 2019). Normally (when $\alpha=1$ and $T=1$), we expect the above inequality to be true. Figure 2 (bottom) confirms that the G-term has an edge over the H-term within a neighborhood around the unaltered model (orange bar), which also corresponds to the Goldilocks zone.
Therefore, it is left to understand when and why Equation 6 is violated. Figure 2 reveals that this occurs at sufficiently low or sufficiently high values of $\alpha$ given a fixed temperature $T=1$. For $\alpha>1$, the norm of the G-term plummets to zero when the logit variance becomes sufficiently large for the softmax output $p^{\prime}$ to collapse to a one-hot distribution for all training samples. In this scenario, $\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}$ is identically zero, as has been observed by Cohen et al. (2021) in the context of progressive sharpening. For us, this implies that $G^{\prime}_{*}(p^{\prime})=J^{\prime\top}[\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}]J^{\prime}$ is identically zero as well, making the Hessian equal to the H-term. Fort & Ganguli (2019) observed the causal relationship between the increased logit variance and vanishing positive curvature but never explained it analytically.
@@
For $\alpha<1$, on the other hand, the G-term does not completely vanish; in fact, it achieves the highest positive curvature at low initialization scales (see the top plots in Figure 2), suggesting that higher entropy predictions should generally be associated with larger positive curvature. According to Equation 5, $\lVert G^{\prime}_{*}\rVert_{2}=\mathcal{O}(\alpha^{2L-2})$ while $\lVert H^{\prime}_{*}\rVert_{2}=\mathcal{O}(\alpha^{L-2})$. Thus, the G-term simply decays faster than the H-term as $\alpha\rightarrow 0$, eventually letting it dominate in the Gauss-Newton decomposition. The higher decay rate of the G-term comes from the cross-class product of logit gradients that vanish as $\mathcal{O}(\alpha^{L-1})$ each. Thus, we call this phenomenon vanishing logit gradient and emphasize that vanishing loss gradient can in fact co-occur with high positive curvature as we demonstrate in the next section.

### v2

@@ Abstract
The second-order properties of the training loss have a massive impact on the optimization dynamics of deep learning models. Fort & Scherlis (2019) discovered that a large excess of positive curvature and local convexity of the loss Hessian is associated with highly trainable initial points located in a region coined the “Goldilocks zone”. Only a handful of subsequent studies touched upon this relationship, so it remains largely unexplained. In this paper, we present a rigorous and comprehensive analysis of the Goldilocks zone for homogeneous neural networks. In particular, we derive the fundamental condition resulting in excess of positive curvature of the loss, explaining and refining its conventionally accepted connection to the initialization norm. Further, we relate the excess of positive curvature to model confidence, low initial loss, and a previously unknown type of vanishing cross-entropy loss gradient. To understand the importance of excessive positive curvature for trainability of deep networks, we optimize fully-connected and convolutional architectures outside the Goldilocks zone and analyze the emergent behaviors. We find that strong model performance is not perfectly aligned with the Goldilocks zone, calling for further research into this relationship.
@@ 1 Introduction
Figure 1: The Goldilocks zone is an area of excess of positive curvature of the loss. Left: Originally, the Goldilocks zone is observed for a narrow range of initialization scales. Middle: Setting the appropriate softmax temperature $T$ allows for excess positive curvature at initialization of any norm. Right: Recreating the Goldilocks zone at an unscaled initialization just by varying $T$.
@@
While the Hessian is extensively studied throughout training and at convergence, fewer works focus on the initialization stage. Recently, Fort & Scherlis (2019) discovered the Goldilocks zone—a region of the optimization space marked by excessive positive curvature $\text{Tr}(H)/\lVert H\rVert_{F}$ and local convexity of the loss Hessian $H$ (see Figure 1). The anomalous readings of these metrics are recorded at a certain distance from the origin of the configuration space where some widely used initialization schemes such as Xavier (Glorot & Bengio, 2010) and Kaiming (He et al., 2015) are found. Thus, the Goldilocks zone is believed to be a hollow centered spherical shell that contains a high density of suitable initial points. Indeed, an appropriate parameter norm is crucial to avoid exploding and vanishing signals, and it seems reasonable that initializations of extreme norm (outside the Goldilocks zone) might suffer from this notorious issue. However, internal covariate shift is largely solved in practice by BatchNorm (Ioffe & Szegedy, 2015), and it can be directly accounted for in some special cases. For example, given an $L$-layer homogeneous network $f_{\alpha\theta}$ with an $\alpha$-scaled initialization (Dinh et al., 2017), the appropriate logit variance and gradient norm can be restored by applying a carefully selected softmax temperature $T=\alpha^{L}$ and learning rate $\eta=\mathcal{O}(\alpha^{2})$. These adjustments ensure that $f_{\alpha\theta}$ follows exactly the same training trajectory and has the same initial excess of positive curvature as $f_{\theta}$ (Figure 1). Thus, contrary to the original claims by Fort & Scherlis (2019), the Goldilocks zone cannot be characterized by the initialization norm alone and is not even a subset of the configuration space. Instead, we find a more fundamental condition governing the excess of positive curvature of the loss in Section 3.
@@
Previous works also attempted to describe the Goldilocks zone more formally. Fort & Ganguli (2019) develop an abstraction, the random logit model, in which they show that prevalence of positive curvature of cross-entropy loss vanishes with increased logit variance. Cohen et al. (2021) attribute the decrease in cross-entropy sharpness during optimization to the collapse of the softmax distribution, which also has implications for the Goldilocks zone. However, the phenomenon of excessive positive curvature—its relation to the initialization norm and trainability—remains unexplained. In this work, focusing on homogeneous networks, we derive a comprehensive description of the Goldilocks zone from the Gauss-Newton decomposition (Section 3) of the cross-entropy loss Hessian and formally associate excess of positive curvature with certain properties of the network (Section 4). In particular, we show that near-zero curvature results from a relatively low spectral norm of the G-term in the Hessian decomposition due to either saturated softmax or vanishing logit gradients, which naturally arise for high- and low-norm initializations, respectively. Inside the Goldilocks zone, we prove that the highest excess of positive curvature is observed for networks with low confidence, which in turn is associated with low initial loss and, for balanced datasets, with vanishing expected loss gradient.
@@
When the conditions governing positive curvature of the loss are established, we inquire about their relation to model trainability and optimization dynamics. Fort & Scherlis (2019) find that optimization on random subspaces of the parameter space converges only if they intersect the Goldilocks zone. For unconstrained optimization, Gur-Ari et al. (2018) discovered that gradients are mainly confined to a low-rank subspace spanned by the Hessian top eigenvectors and vanish along flatter directions. Since the excess of positive curvature manifests a larger separation between bulk and outlier eigenvalues, it should then be associated with a more robust top-eigenspace and, intuitively, a more informative training signal. Motivated to unveil this relationship, we use gradient descent to optimize homogeneous networks using a wide spectrum of initialization norms and learning rates and taxonomize the emergent behaviors. Interestingly, we find that successful training is not necessarily well aligned with the Goldilocks zone. We demonstrate setups where the slightest increase in the initialization norm of LeNet-5 leads to degenerate learning despite happening well within boundaries of the Goldilocks zone. These dynamics are marked by an increasing amount of zero logits and, to the best of our knowledge, we are the first to report this behavior.
@@
This paper conducts an extensive study of the Goldilocks zone of homogeneous neural networks, both analytically and empirically. Our code is available at https://GitHub.com/avysogorets/goldilocks-zone.
  • In Section 3, we demonstrate that the Goldilocks zone is not characterized by the initialization norm alone, refining prior beliefs of Fort & Scherlis (2019). Instead, we derive a more fundamental condition resulting in excess of positive curvature and find that it disappears due to saturated softmax on one end and vanishing logit gradients on the other.
  • In Section 4, we closely study the interior of the Goldilocks zone and analytically associate excess of positive curvature with low model confidence, low initial loss, and low cross-entropy gradient norm.
  • In Section 5, we report the evolution and performance of scaled homogeneous networks when optimized by gradient descent both inside and outside the Goldilocks zone. Our investigation shows that excess of positive curvature is an imperfect estimator of the initialization trainability and exhibits a range of interesting effects for initializations on the edge.
@@ 2 Preliminaries & Notation
By default, $T=1$, in which case we refer to softmax simply as $\sigma$. The corresponding cross-entropy loss is $\ell(p,y)=-\log p_{y}$. The Hessian matrix $H\in\mathbb{R}^{P\times P}$ holds the second derivatives of the loss at $\theta$: $H_{ij}=\partial^{2}\ell/\partial\theta_{i}\partial\theta_{j}$. In principle, all of the above quantities depend on one particular or a batch of inputs $(X^{\mu},y^{\mu})$; we use the superscript $\mu$ to make this dependence explicit where needed. Since we are often concerned with the magnitude of network outputs at different parameter scales, we assume the inputs are standardized and bounded.
Homogeneous networks.
We restrict our analysis to homogeneous models satisfying $f_{\alpha\theta}=\alpha^{L}f_{\theta}$ for any scalar $\alpha>0$ where $L$ is the number of layers in $f$. This is a rather technical assumption: homogeneous models are widely used in practice and include ReLU networks without biases. Fort & Scherlis (2019) focus on initially homogeneous models by setting biases to zero. Inhomogeneous networks encompass a wide range of architectures and design choices that require case-by-case analysis depending on the source of inhomogeneity (e.g., non-zero biases, residual connections, inhomogeneous activation functions, etc.). Many inhomogeneous cases may be quite degenerate in the context of initialization scaling. For example, upscaling the initialization of TanH-networks simply saturates the activations, blocking any signal propagation. Downscaling these networks sufficiently hard makes them into linear models. For another example, consider ReLU networks with non-zero biases and at least one hidden layer, which are inhomogeneous due to the different scales of biases and activations in hidden layers. When $\alpha\ll 1$, the output signal is dominated by biases; when $\alpha\gg 1$ biases vanish in magnitude compared to the corresponding activations, making the model almost homogeneous. In Section 6, we revisit inhomogeneous models and suggest how our analyses can be extended to these architectures as well.
Excess of positive curvature.
@@
In their work, Fort & Scherlis (2019) define the Goldilocks zone rather informally as a region of excessive positive curvature of the loss, which can be unmistakably identified in Figure 1. of positive curvature of the loss is defined as the excess of positive eigenvalues $\{\lambda_{i}\}_{i=1}^{P}$ of a loss Hessian $H$:
$\frac{\text{Tr}(H)}{\lVert H\rVert_{F}}=\frac{\sum_{i=1}^{P}\lambda_{i}}{\sqrt{\sum_{i=1}^{P}\lambda_{i}^{2}}}.$ (2)
@@
Throughout the remainder of the paper, we refer to this quantity simply as positive curvature for fluid presentation. Another metric used by Fort & Scherlis (2019) is local convexity of the loss, which refers to the fraction of positive eigenvalues of $H$. These two metrics are intimately related and can detect the Goldilocks zone independently of each other, so we focus only on positive curvature throughout this paper. Since computing the full Hessian is intractable for most modern architectures, Fort & Scherlis (2019) substitute it with a projection $H_{d}=R^{T}HR$ onto a low-rank random subspace with basis defined by a sparse matrix $R\in\mathbb{R}^{P\times d}$ with orthogonal columns. Given an initialization $\theta$, this amounts to computing the Hessian of the training loss of the model parameterized by $R\hat{\theta}+\theta$ with respect to latent $d$-dimensional parameters $\hat{\theta}$ at the origin of the chosen low-rank space, which is much more accessible. We adopt the same strategy and assume that the first- and second-order derivatives with respect to $\theta$ more generally represent derivatives with respect to trainable $d$-dimensional parameters, which can be latent parameters $\hat{\theta}$ or the original model weights if we let $d=P$ and $R=I_{P}$. This technical nuisance affects none of our analyses, but we add further comments as it becomes necessary. We verify the validity of this practical approach by comparing it to Hutchinson’s stochastic trace estimation in Appendix F (Hutchinson, 1990).
@@ 3 Revisiting the Goldilocks Zone
Fort & Scherlis (2019) introduced the Goldilocks zone as a region of the parameter space with an excess of positive curvature and local convexity of the loss function, as defined in Section 2. Starting from a gold standard Kaiming initialization $\theta_{0}$, Fort & Scherlis (2019) record these statistics over a ray $\{\alpha\theta_{0}\colon\alpha>0\}$ and read unusually high values when $\alpha$ falls within a relatively narrow range centered around $\alpha=1$ (Figure 1 left). Moreover, they find that SGD constrained to a random subspace is successful only if it intersects the Goldilocks zone. Based on these observations, the authors suggest that the Goldilocks zone is a thick, hollow spherical shell about the origin in the configuration space, which is densely populated with initial points amenable for training.
A simple example shows that, strictly speaking, this visually appealing representation needs revision. For homogeneous models, the $\alpha$-scale transformation $\theta\rightarrow\alpha\theta$ with $\alpha>0$ effectively scales the underlying configuration space by $\alpha$ (Dinh et al., 2017). We shall derive next that the cross-entropy loss landscape scales together with the configuration space when the softmax temperature satisfies $T=\alpha^{L}$, restoring positive curvature of the loss of the scaled model $f_{\alpha\theta}$ to that of the original model $f_{\theta}$ (see the middle plot in Figure 1). Thus, we argue that initialization norm has a coincidental relationship to the Goldilocks zone, calling for a refined, analytically driven characterization.
@@
$H_{ij}=\underbrace{\sum_{k=1}^{K}\sum_{c=1}^{K}\frac{\partial z_{k}}{\partial\theta_{i}}\left[\nabla^{2}_{z}\ell\right]_{kc}\frac{\partial z_{c}}{\partial\theta_{j}}}_{\text{G-term }(\mathcal{G})}+\underbrace{\sum_{k=1}^{K}\left[\nabla_{z}\ell\right]_{k}\frac{\partial^{2}z_{k}}{\partial\theta_{j}\partial\theta_{i}}}_{\text{H-term }(\mathcal{H})}\,,$ (4)
where $\nabla^{2}_{z}\ell=\text{diag}(p)-pp^{\top}$ is the Hessian of the loss with respect to the logits and $\nabla_{z}\ell=p-\text{OH}(y)$ where $p$ is the softmax output and $\text{OH}(y)$ is a one-hot encoded tagret (Cohen et al., 2021). To emphasize the dependence of the G-term and H-term on $p$ we sometimes refer to them as $\mathcal{G}(p)$ and $\mathcal{H}(p)$, respectively. Note that $\mathcal{G}(p)=J^{\top}[\text{diag}(p)-pp^{\top}]J$ where $J$ is the Jacobian matrix. For the $\alpha$-scaled model $f^{\prime}$ used together with softmax $\sigma_{T}$, we similarly get $\nabla^{2}_{z^{\prime}}\ell^{\prime}=T^{-2}\left[\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}\right]$, $\nabla_{z^{\prime}}\ell^{\prime}=T^{-1}(p^{\prime}-\text{OH}(y))$. Combining this with homogeneity of gradients, we derive the Gauss-Newton decomposition of the Hessian of $f^{\prime}$:
$H^{\prime}$ $=\mathcal{G}^{\prime}+\mathcal{H}^{\prime}=\frac{\alpha^{2L-2}}{T^{2}}\mathcal{G}(p^{\prime})+\frac{\alpha^{L-2}}{T}\mathcal{H}(p^{\prime}).$ (5)
Linearity of differentiation ensures that this equation is valid for the full-batch Hessian, too. Indeed, the full G-term and the full H-term are just the averages of the per-sample quantities $\mathcal{G}^{\mu}$ and $\mathcal{H}^{\mu}$ defined in Equation 4, respectively. Thus, we will abuse this notation and refer to their full counterparts in the same way where appropriate. In this case, the probability vectors $p$ and $p^{\prime}$ can be viewed as matrices.
Returning to our discussion on the shape of the Goldilocks zone, we remark that $T=\alpha^{L}$ yields $H^{\prime}=\alpha^{-2}H$ (as this temperature ensures $p=p^{\prime}$). Since the measures of positive curvature and local convexity are robust to scaling of the Hessian matrix, we conclude that initialization of any norm can be in the Goldilocks zone provided an appropriate softmax temperature (see Figure 1). Now that the connection between the initialization norm and the Goldilocks zone is much less credible, we leverage Equation 5 to establish the fundamental principles governing positive curvature in neural networks.
@@
The Gauss-Newton decomposition in Equation 4 is a common entry point for many studies on the Hessian of large neural networks. The Hessian exhibits a “bulk-outlier” eigenspectrum with the majority of eigenvalues small and clustered around zero and only a handful of large positive outliers (Sagun et al., 2016; Gur-Ari et al., 2018; Ghorbani et al., 2019). This decomposition is inherited from the individual spectra of $\mathcal{G}$ and $\mathcal{H}$ with the top and the bulk eigenvalues attributed to these two terms, respectively. We present a survey of works concerned with this phenomenon in Appendix A.
@@
We are now in position to describe the exact conditions that result in prevalence of positive curvature—the hallmark of the Goldilocks zone. The eigenstructures of the individual terms in the Gauss-Newton decomposition of the Hessian suggest that the G-term is the one and only source of positive curvature of the loss, and so
$\lVert\mathcal{G}^{\prime}\rVert_{2}\gtrsim\lVert\mathcal{H}^{\prime}\rVert_{2}$ (6)
is a necessary and sufficient condition for the Goldilocks zone. We observe this exact correspondence in Figure 2. Otherwise, when $\lVert\mathcal{G}^{\prime}\rVert_{2}$ is sufficiently small relative to $\lVert\mathcal{H}^{\prime}\rVert_{2}$, the bulk-like eigenspectrum of the H-term swallows the outliers of the G-term and is inherited by the Hessian, resulting in a near-zero positive curvature. This sudden change in the eigenstructure is known as the BBP phase transition (Baik et al., 2005; Fort & Ganguli, 2019). Normally (when $\alpha=1$ and $T=1$), we expect the above inequality to be true. Figure 2 (bottom) confirms that the G-term has an edge over the H-term within a neighborhood around the unaltered model (orange bar), which also corresponds to the Goldilocks zone.
@@
Therefore, it is left to understand when and why Equation 6 is violated. Figure 2 reveals that this occurs at sufficiently low or sufficiently high values of $\alpha$ given a fixed temperature $T=1$. For $\alpha>1$, the norm of the G-term plummets to zero when the logit variance becomes sufficiently large for the softmax output $p^{\prime}$ to collapse to a one-hot distribution for all training samples. In this scenario, $\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}$ is identically zero, as has been observed by Cohen et al. (2021) in the context of progressive sharpening. For us, this implies that $\mathcal{G}^{\prime}(p^{\prime})=J^{\prime\top}[\text{diag}(p^{\prime})-p^{\prime}p^{\prime\top}]J^{\prime}$ is identically zero as well, making the Hessian equal to the H-term. Fort & Ganguli (2019) observed the causal relationship between the increased logit variance and vanishing positive curvature but never explained it analytically.
For $\alpha<1$, on the other hand, the G-term does not completely vanish; in fact, it achieves the largest excess of positive eigenvalues at low initialization scales (see the top plots in Figure 2), suggesting that higher entropy predictions should generally be associated with larger positive curvature. According to Equation 5, $\lVert\mathcal{G}^{\prime}\rVert_{2}=\mathcal{O}(\alpha^{2L-2})$ while $\lVert\mathcal{H}^{\prime}\rVert_{2}=\mathcal{O}(\alpha^{L-2})$. Thus, the G-term simply decays faster than the H-term as $\alpha\rightarrow 0$, eventually letting it dominate in the Gauss-Newton decomposition. The higher decay rate of the G-term comes from the cross-class product of logit gradients that vanish as $\mathcal{O}(\alpha^{L-1})$ each. Thus, we call this phenomenon vanishing logit gradient and emphasize that vanishing loss gradient can in fact co-occur with high positive curvature as we demonstrate in the next section.